OpenStudy (anonymous):

The area of a rectangular piece is 280 if the length is 5 meters less and the width was 1 meter more the piece would be a square Write and equation to find the width Find the width

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Ashleyisakitty @TheSmartOne

TheSmartOne (thesmartone):

Welcome to OpenStudy @Jgerftnyt What have you understood from your question?

OpenStudy (anonymous):

That I have to make an equation and I'm confused from the steps could you please help me

TheSmartOne (thesmartone):

What do you make of this sentence: if the length is 5 meters less and the width was 1 meter more the piece would be a square

OpenStudy (anonymous):

Y-5 and x+1

OpenStudy (anonymous):

How do I make the equation

TheSmartOne (thesmartone):

and it says it would make a square, and a square has sides that are all equal, right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What's next???

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

L - 5 = W + 1 Solve for L

TheSmartOne (thesmartone):

and then you plug that into \(\sf L \times W = 280\) and you set it equal to 0 and you'll have a quadratic equation that you'll need to solve.

OpenStudy (anonymous):

@TheSmartOne so x+1 =y-5. X+6=y whats next

TheSmartOne (thesmartone):

lets use L and W because they stand for length and width and it will make it easier for us :)

TheSmartOne (thesmartone):

so we have L = W+6 Plug that into \(\sf\Large L \times W = 280\) so we get \(\sf\Large (w+6)w=280\)